Find all possible rational canonical form for a 6×6 matrix over R with minimal polynomial m(t)= (t2 -2t +3)(t+1)2

1
Expert's answer
2021-01-29T00:33:06-0500

Find comparision matrices of invariant factors:

for "t+1" :

"C_1=[-1]"

for "t^2-2t+3" :

"C_2=\begin{pmatrix}
 0 & -3 \\
 1 & 2
\end{pmatrix}"

for "m(t)= (t^2 -2t +3)(t+1)^2=t^4-t^2+t+3"

"C_4=\begin{pmatrix}
 0 & 0&0&-3 \\
 1 & 0&0&-1\\
 0&1&0&1\\
 0&0&1&0 \\
\end{pmatrix}"


for "(t^2 -2t +3)(t+1)=t^3-t^2+t+3"

"C_3=\begin{pmatrix}
 0 & 0&-3 \\
 1 & 0&-1\\
 0&1& 1 \\
\end{pmatrix}"


If "m(t)=q^{m_1}_1...q^{m_i}_i" , then

size of block: "n_j(deg(q_{0j}))"

# blocks: "\frac{dim (Null(q_{0j}(A)))}{deg(q_j)}"


Possible rational canonical form for a 6×6 matrix: